Part C 31.5 g of NaOH in 370. mL of a NaOH solution Express your answer with the appropriate units. Molarity ValueUnits Submit Provide Feedback
Answer #1

Molarity = moles of solute / Liters of solution

converting 31.5 g NaOH into moles: 31.5 g /40.0 g/mol = 0.7875 moles of solute

converting 370 ml into Liters: 370 mL x ( 1 L /1000 mL ) = 0.370 L solution

Molarity = 0.7875 / 0.370 = 2.13 mol/L ( 1 mol/L = 1 M )

Thank you